A massless, frictionless pulley has a rope with two weights suspended from it. The masses of the blocks are as follows: MA = 2m and MB = m. What is the tension in the rope when the blocks are released from rest?

a. 2 mg b. 3/2 mg c. 4/3 mg d. 1/2 mgA massless, frictionless pulley has a rope with tw
Answer #1

a1 a2 Mb Ma

From an above diagram, we have

down positive : mb g - T = mb a

T = (mb g - mb a)

up positive : T - ma g = ma a

(mb g - mb a) - ma g = ma a

(mb g - ma g) = (mb a + ma a)

a = (mb - ma) g / (mb + ma)

using an equation, we have

T = (mb g - mb a)

T = mb g - mb [(mb - ma) g / (mb + ma)]

T = mb g [(mb + ma) g / (mb + ma)] - mb g [(mb - ma) g / (mb + ma)]

T = [mb (mb + ma) - mb (mb - ma) / (mb + ma)] g

T = [2 mb ma / (mb + ma)] g

where, mb = mass of block B = m

ma = mass of block A = 2m

then, w get

T = 2 [(m) (2m) / (m) + (2m)] g

T = (4 m2 / 3m) g

T = (4/3) mg